please help me in the trigonometry identities and equations

please help  me in the trigonometry identities and equations
 

Grade:11

2 Answers

Arun
5 years ago
Dear Sara
please ask question here. You can also attach an image. I will be happy to help you.
 
Regards
Arun (askIITians forum expert)
SARA
5 years ago
thank you vewry much .I need to solve basic trigonometry equations.sin2x=0
. tan(2x+128)=0
cot(3x-1)=-1
2cos2x-1=0
tan(2x-5)-cot(x+1)=0